OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P?

OpenStudy (lgbasallote):

equation of a circle is \[(x-h)^2 + (y-k)^2 = r^2\] where (h,k) is the center and r is the raidus does that help

OpenStudy (anonymous):

(x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 64 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 8

OpenStudy (lgbasallote):

so does your flaunting of the choices mean you didnt get what i said?

OpenStudy (anonymous):

Yeah. Haha because none of them matched the - + - pattern

OpenStudy (lgbasallote):

- - means +

OpenStudy (lgbasallote):

^does that make sense?

OpenStudy (anonymous):

Yes I know that.

OpenStudy (lgbasallote):

so what's your problem then?

OpenStudy (lgbasallote):

maybe if i show you a demo...

OpenStudy (anonymous):

I have to choose which of those defines P.

OpenStudy (lgbasallote):

the center of a circle is (3, -4) and radius is 5 he equation would be \[(x- 3)^2 + (y - -4)^2 = 5^2\] \[\implies (x-3)^2 + (y+4)^2 = 25\] got the idea now?

OpenStudy (anonymous):

OH I think so, thanks :)

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (anonymous):

The solution is: (x-2)^2+(y+1)^2=64